OpenStudy (anonymous):

Find the next three terms in the geometric sequence. 4, –8, 16, –32, ...

OpenStudy (anonymous):

the next term = previous term x -2 or -32 x -2 = 64 64 x -2 = -128 128 x -2 = 256

OpenStudy (anonymous):

what i see here is a pattern of multiplying by -2 4 x -2 = -8 -8 x -2 = 16 16 x -2 = -32 -32 x -2 = 64 64 x -2 = -128 -128 x -2 = 256

OpenStudy (anonymous):

x as in times (*) not (x)

OpenStudy (anonymous):

thanks guys!!!